OpenStudy (anonymous):

how do you find a^n for the following geometric sequence a^1=6,r=-2

OpenStudy (anonymous):

Howz the sequence goes ?

OpenStudy (amistre64):

a(n+1) = a0 * -2n i think

OpenStudy (amistre64):

hard to tell where to start; some suggest a0 some suggest a1

OpenStudy (anonymous):

yeah,the question doesn't gives all information.

OpenStudy (amistre64):

it gives all the pertenant info, but the method seems to be more material specific

OpenStudy (amistre64):

if its starts at a1 then we have: an = 6*-2(n-1) a1 = 6*-2(0) = 6

OpenStudy (amistre64):

if it starts at a0 then: an = -3*-2n

OpenStudy (anonymous):

I am really having trouble understanding the problem,what is \[a^n\] means here? Is the sequence like \[ a,a^2,a^3,a^4 \cdots \] ?

OpenStudy (amistre64):

i think its a typo for a_n

OpenStudy (amistre64):

\[a_n=a_1*-2(n-1)\]

OpenStudy (anonymous):

But when \[ a_1=1 \] (first term) and r=common ratio r=-2,shouldn't \[a_n = a_1 \times r^{(n-1)}=6 \times (-2)^{(n-1)} \] ...

OpenStudy (anonymous):

and how is this \[ a_n=a_1*-2(n-1) \] holds? What exactly I am missing here ?:/

OpenStudy (amistre64):

geometric, if i recall is a multiplier between terms; the functional notation results in an exponential result tho

OpenStudy (amistre64):

if we are just going from term to term tho, its a multiplier;

OpenStudy (anonymous):

Sorry but I simply can't understand your argument .. I am talking about http://en.wikipedia.org/wiki/Geometric_progression and you ? :)

OpenStudy (amistre64):

"Such a geometric sequence also follows the recursive relation"

OpenStudy (amistre64):

i might be messing it up tho, but that recursive relation is what i was considering

OpenStudy (anonymous):

yep,and that relation is \[ a_n =r \times a_{(n-1)} \] but how does that explain \[ a_n=a_1*-2(n-1) \] ????!!!

OpenStudy (amistre64):

\[a_1 = 6\] \[a_2 = 6(-2)=-12\] \[a_3 = 6(-2)(-2)=24\] \[a_3 = 6(-2)(-2)(-2)=-48\] \[a_n = a_1*(-2)^{n-1}\] i see what i did, i forgot to put it as an exponent

OpenStudy (anonymous):

Aha,thats makes sense now :)

OpenStudy (amistre64):

im old and feeble, thats my excuse lol

OpenStudy (anonymous):

Time to change your diapers I guess ?:P

OpenStudy (amistre64):

still got some load to them, maybe in an hour ;)

OpenStudy (anonymous):

lolz :D Cheers ! :)